# A measure-valued HJB perspective on Bayesian optimal adaptive control

### v1

@@ Abstract.
We consider a Bayesian adaptive optimal stochastic control problem where a hidden static signal has a non-separable influence on the drift of a noisy observation. Being allowed to control the specific form of this dependence, we aim at optimising a cost functional depending on the posterior distribution of the hidden signal. Expressing the dynamics of this posterior distribution in the observation filtration, we embed our problem into a genuinely infinite-dimensional stochastic control problem featuring so-called measure-valued martingales. We address this problem by use of viscosity theory and approximation arguments. Specifically, we show equivalence to a corresponding weak formulation, characterise the optimal value of the problem in terms of the unique continuous viscosity solution of an associated HJB equation, and construct a piecewise constant and arbitrarily-close-to-optimal control to our main problem of study.
@@ 1. Introduction
We consider a Bayesian adaptive optimal stochastic control problem in continuous time on an infinite horizon. There is an underlying static signal which cannot be observed directly but only via a noisy observation. More precisely, the drift of this observation process has a non-linear dependence on the signal and we are allowed to control the specific form of this dependence. We take a Bayesian view on the estimation problem and assume that the prior is known and subsequently update our beliefs. The aim is to optimise an objective depending on both the control and the posterior distribution of the hidden signal. The typical interpretation is that we are searching for the hidden ‘signal’ and based on what we’ve learned so far we may choose different ‘search actions’ which might be more or less effective but at the same time more or less costly.
In mathematical terms, the observation process is given by
$\mathrm{d}Y^{u}_{t}=h(u_{t},X)\mathrm{d}t+\mathrm{d}W_{t},\quad Y^{u}_{0}=0,$
where $X$ is an unobservable random variable, $u$ is the chosen control, and the function $h$ specifies the spatial dependence on the signal corresponding to different actions. For any given control, the observation process generates a filtration, say $\mathcal{Y}^{u}$, which is explicitly control dependent. The posterior distribution, also referred to as the filter, is then given by
@@
Given a discount rate $\beta$ and a running cost function, say $k$, depending on the control and the current state of the posterior distribution, the aim is to optimise an objective of the form
$\mathbb{E}\int_{0}^{\infty}e^{-\beta t}k(\pi^{u}_{t},u_{t})\mathrm{d}t;$
@@
the optimisation is over a class of controls $u$ which are adapted to their own generated filtration $\mathcal{Y}^{u}$, that is, controls that rely only on the information on $X$ generated from observing $Y^{u}$.
@@
We address this problem by use of dynamic programming arguments and stochastic control methods. To this end, we first express the dynamics of the posterior distribution in the observation filtration; the filter can then be viewed as a controlled measure-valued martingale whose dynamics is driven by a Brownian motion. We then introduce two thereto related formulations: A weak problem formulation and an approximate formulation where one restricts to piecewise constant controls. A priori, it is not clear that either of these formulations coincide. However, our main result establishes that the value of our original problem, the value of the weak formulation, and the limiting value of the approximate formulation as the time grid is refined, are equal in value. A posteriori, it is thus clear that the additional information available within the weak formulation does not alter the value of the problem. The proof bears resemblance to the approximation approach introduced in [4] as well as the stochastic Perron method developed in [6]; specifically, relying on the viscosity theory for MVMs developed in [18], it proceeds by establishing that the limiting and weak value functions are sub and super solutions, respectively, of an associated HJB equation to which appropriate comparison results applies. In particular, we obtain that the value of the problem can be characterised in terms of the unique continuous viscosity solution of this HJB equation. Finally, having constructed an optimiser for a piecewise constant problem with a fine enough time grid, returning to our original problem and relying on uniqueness in law for the controlled filtering equation, we construct a piecewise constant and arbitrarily-close-to-optimal strategy for our main problem of interest.
Our main problem of study is notably given in a strong formulation where the observation filtration has an explicit dependence on the control. At the same time, admissible controls are constrained to depend on the information provided by the observation process only. When aiming at picking a ‘good control’ there is thus a trade-off present since such controls must balance improving the estimation and minimizing the cost functional. The effect of considering strong formulations within stochastic control problems with unobservable components was recently reviewed in [16]. In particular, they illuminated the link between controlling the information flow and the notions of dual effect (referring to the control’s effect on higher order moments of the posterior distribution) and the much-studied trade-off between exploration and exploitation (referring to the knowledge of the unobservable component contra cost optimisation); we refer to that article for further details and references on the topic. Within the literature on optimal control under partial information, controls that are restricted to depend on the observation process only are often referred to as ‘strict-sense’ controls in contrast to so-called ‘wide-sense’ controls which are allowed to depend on a somewhat larger filtration. The latter were introduced in [25] and used in e.g. [23] and [8] where it was shown, among other things, that the value of their problem remained the same regardless of whether strict- or wide-sense controls were used; our main result is in line with this. In this literature, it is also common to consider a fixed filtered probability space on which admissible controls are defined and then let the controlled dynamics enter the problem via a change of measure. Crucially, the relevant filtration then becomes independent of the chosen control (see e.g. [9] for the case of strict-sense controls and [33] for the case of wide-sense controls). It is also common to optimise over tuples of this form. Our main result verifies that for the problem at hand, such variants of weak formulations would agree in value.
@@
Our work is notably inspired by the contribution of [16]. Following their approach, we introduce our strong problem formulation by defining a class of pre-admissible controls and restricting in turn to the subset of controls which rely on the information provided by the observations only. Our usage of stochastic Perron and Barles-Souganides type arguments for establishing equivalence between our main problem of interest and the weak and approximate formulations thereof, is also closely inspired by their arguments. The key contribution of our work is that we allow for a more general dependence on the signal in the drift coefficient of the observation process; while we restrict to an action space enforcing a certain polynomial structure, we crucially allow for infinite polynomials. While this might seem an innocent extension at first, it has far-reaching consequences: In contrast to the problem studied in [16], our problem can not be reduced to a finite-dimensional one. This calls for a genuinely infinite-dimensional analysis working directly with the probability-measure–valued filters. We carry out such a study by relying on the viscosity theory developed for stochastic control problems featuring MVMs in [18]. In particular, we herein develop approximation arguments analogous to the results of Barles-Souganides [4] for this infinite-dimensional set-up; we find those developments to be of independent interest and hope that they will be useful also in other contexts. Apart from extending the existing theory on Bayesian optimal control problems, we thus also provide a concrete example of how the viscosity theory of [18] can be applied; in particular, for the first time, we establish continuity of the value function for this genuinely infinite-dimensional set-up. Our analysis also includes an original result on the stability of controlled MVMs in terms of their initial condition, a result which we also find to be of independent interest.
We now discuss the related literature. Within a more applied context, Bayesian search problems have been studied in e.g. [1, 20, 27, 44]. This literature aside, there are essentially three streams of research to which our problem relates: First, our problem bears resemblance to the problem of stochastic control under partial information; the difference being that our signal is kept constant while we control the observation rather than the signal itself. For this class of problems, the existence of wide-sense optimal controls were studied in [24], [25] and [23]; see also [14] and [29]. The predominant approach in this literature has been to address the original problem via a reformulation of the objective enabling viewing it as a control problem where the conditional distribution of the hidden signal plays the role of the controlled state variable. The approach taken herein is of course closely related to this stream of research and we will review the relevant literature in more detail in Section 2. Second, motivated by aims similar to ours, closely related problems have been considered within an optimal stopping framework. For example, in [21], the authors let an unobservable state influence the drift of the state process and their aim is to estimate this drift as accurately but also as quickly as possible. Considering a Bayesian framework they formulate it as an optimal stopping problem which they subsequently address. An extended version of such a stopping problem was also studied in [22]. We refer to those articles for further references. Third, our formulation notably includes the problem of minimising, over admissible controls, an objective of the form [^1] Notably there is still a trade-off present when choosing the control since it affects both the cost and the estimation of the unknown signal; as before, this is a consequence of the fact that we consider a strong formulation where the choice of the control influences the available information through the filtration.
(1) $\mathbb{E}\int_{0}^{\infty}e^{-\beta t}\tilde{k}(X,u_{t})\mathrm{d}t,$
where $\tilde{k}$ is some given cost function; this can be seen by defining $k(\mu,v)=\mu(\tilde{k}(\cdot,v))$, for any probability measure $\mu$ and action $v$. Closely related problems have been studied previously in the literature:
  • In [8], the authors consider a problem where the observation process follows a dynamics corresponding to letting $h(v,x)=vx$, for real-valued controls, and where the aim is to optimise an objective similar to (1) but with a running cost depending on the observation process itself. Specifically, they consider a cost of the form $\tilde{k}(Y^{u}_{t})$ for $\tilde{k}(y)=y^{2}$, $y\in\mathbb{R}$. In [33], their results were extended to the case of $\tilde{k}$ being an arbitrary even, convex function of exponential growth. The corresponding finite-horizon problem was studied in [34]. Those articles obtained explicit results but under rather restrictive assumptions on the prior.
  • More recently, the above-mentioned article [16], considered an observation process whose dynamics is again similar to ours but with $h(u_{t},X)$ replaced by $b(t,Y^{u}_{t},u_{t})X$, for some function $b$, and with a similar controlled volatility coefficient appearing in front of the noise term; they also allowed for a multi-dimensional signal. They considered the problem of optimising a finite-horizon objective with a running cost of the form $\tilde{k}(t,Y^{u}_{t},X,u_{t})$ and a terminal cost of analogous form.
While those papers motivated our work, we note that our set-up is different: while we crucially allow for a non-linear dependence on the signal in the drift of the observation process as well as a non-linear dependence on the posterior distribution in the cost functional, neither the drift nor the cost has any explicit dependence on the observation process in our set-up. In order to render the presentation focused on the difficulties arising from the non-linear dependence on the signal, we choose to develop the theory for the case without such an explicit dependence on the observation process. We believe, however, that our results could be extended to cover also this situation but leave it for future research to thoroughly investigate this.
@@
The remainder of the article is organized as follows: In Section 2 we introduce our main problem of interest, express the dynamics of the involved processes in the observation filtration, and discuss the related literature; in Section 3 we introduce our weak problem formulation as well as the piecewise constant approximate formulation and provide our main result; Section 4 presents our stability results and Section 5 is devoted to the proof of our main result.
@@
Throughout, $\mu_{0}$ will be a given fixed probability measure on $\mathbb{R}$. We denote by $\mathcal{P}$ the space of probability measures on $\mathbb{R}$ whose support is contained in the support of $\mu_{0}$; we equip this space with the topology of weak convergence rendering it a Polish space. Since we will assume that $\mu_{0}$ has compact support, we can work with e.g. the Wasserstein-$1$ metric, which we denote by $\mathcal{W}$. That $\mu_{0}$ has compact support also implies that the space $\mathcal{P}$ is compact. We denote by $\mathcal{P}^{s}$ the (closed) subset of probability measures supported in one single point. We write $C_{b}$ for the bounded continuous functions on $\mathbb{R}$.

### v2

@@ Abstract.
We consider a Bayesian adaptive optimal stochastic control problem where a hidden static signal has a non-separable influence on the drift of a noisy observation. Being allowed to control the specific form of this dependence, we aim at optimising a cost functional depending on the posterior distribution of the hidden signal. Our setup is in sharp contrast to existing work: we include costs that depend on the full posterior distribution in a form that admits a large class of non-linear relationships. Expressing the dynamics of this posterior distribution in the observation filtration, we embed our problem into a genuinely infinite-dimensional stochastic control problem using measure-valued martingales. We address this problem by use of viscosity theory and approximation arguments. Specifically, we show equivalence to a corresponding weak formulation, characterise the optimal value of the problem in terms of the unique continuous viscosity solution of an associated HJB equation, and construct a piecewise constant and arbitrarily-close-to-optimal control to our main problem of study. As a byproduct of our analysis, we also provide a novel stability result for a class of measure-valued SDEs which we believe is of independent interest.
@@ 1. Introduction
We consider a Bayesian adaptive optimal stochastic control problem in continuous time on an infinite horizon. There is an underlying static signal which cannot be observed directly but only via a noisy observation. More precisely, the drift of this observation process has a non-linear dependence on the signal and we are allowed to control the specific form of this dependence. We take a Bayesian view on the estimation problem and assume that the prior is known and subsequently update our beliefs. The aim is to optimise an objective depending both on the control and the posterior distribution of the hidden signal. The typical interpretation is that we are searching for a hidden ‘signal’ and based on what we’ve learned so far we may choose different ‘search actions’ which might be more or less effective but at the same time more or less costly.
Problem formulation
In mathematical terms, given an unobservable random variable $X$ with law $\mu_{0}$, the observation process is given by
(1) $\mathrm{d}Y^{u}_{t}=h(u_{t},X)\mathrm{d}t+\mathrm{d}W_{t},\quad Y^{u}_{0}=0,$
where $u$ is the chosen control and the function $h$ specifies the spatial dependence on the signal for different actions. For any given control, the observation process generates a filtration, say $\mathcal{Y}^{u}$, which is explicitly control dependent. The posterior distribution, also referred to as the filter, is given by
@@
Given a discount rate $\beta$ and a running cost $k$ depending on the control and the current state of the posterior distribution, the aim is to optimise an objective of the form
(2) $\mathbb{E}\int_{0}^{\infty}e^{-\beta t}k(\pi^{u}_{t},u_{t})\mathrm{d}t;$
@@
the optimisation is performed over controls $u$ which are adapted to their own generated filtration $\mathcal{Y}^{u}$, that is, controls that rely only on the information on $X$ generated from observing $Y^{u}$.
Results and methodology
We address the above problem by use of dynamic programming arguments and stochastic control methods. To this end, we first express the dynamics of the posterior distribution in the observation filtration; specifically, for any bounded continuous function $\varphi$, writing $\pi^{u}_{t}(\varphi)$ for $\int\varphi(x)\pi^{u}_{t}(\mathrm{d}x)$ etc., we have that
(3) $\mathrm{d}\pi^{u}_{t}(\varphi)=\big[\pi^{u}_{t}\big(\varphi(\cdot)h(u_{t},\cdot)\big)-\pi^{u}_{t}\big(\varphi\big)\pi^{u}_{t}\big(h(u_{t},\cdot)\big)\big]\mathrm{d}I^{u}_{t},\;\;\pi^{u}_{0}=\mu_{0},$
@@
where $I^{u}$ is a Brownian motion in the observation filtration. The filter can thus be viewed as a controlled measure-valued martingale (MVM) with dynamics driven by a Brownian motion. We introduce two related formulations: A weak problem formulation where the optimisation is over weak solutions to (3), and an approximate formulation where one restricts to measure-valued processes solving (3) for piecewise constant controls. A priori, it is not clear that either of these formulations coincide. However, our main result establishes that the value of our original problem, the value of the weak formulation, and the limiting value of the approximate formulation as the time grid is refined, are equal in value. A posteriori, it is thus clear that the additional information available within the weak formulation does not alter the value of the problem. We also establish that the value of the problem can be characterised in terms of the unique continuous viscosity solution of an associated HJB equation. Finally, we construct a piecewise constant and arbitrarily-close-to-optimal strategy for our main problem of interest.
@@
Our proof bears resemblance to the Barles-Souganides approximation method introduced in [4], as well as to the stochastic Perron approach developed in [7]; it proceeds by establishing that the limiting and weak value functions are sub and super solutions, respectively, of an associated HJB equation, by sandwiching our main problem of interest in between the two, and by applying appropriate comparison results. Our method is notably inspired by the contributions in [14]. The key feature of our work is that we allow for a more general dependence on the signal in the drift of the observation process; while we restrict to an action space enforcing a certain polynomial structure, we crucially allow for infinite polynomials. In contrast to the problem studied in [14], our problem can therefore not be reduced to a finite-dimensional one but calls for a genuinely infinite-dimensional analysis where one works directly with the probability-measure–valued filters. We carry out such an analysis by relying on the viscosity theory developed for stochastic control problems with MVMs in [16]. In particular, we develop Barles-Souganides type approximation arguments for this infinite-dimensional setup; we find those developments to be of independent interest and hope that they will become useful also in other contexts. In order to establish continuity of the value function, we also provide a result on the stability of controlled MVMs in terms of their initial condition; we find this result too to be of independent interest.
Our main problem of study is notably given in a strong formulation where the observation filtration has an explicit dependence on the control. At the same time, admissible controls are constrained to depend on the information provided by the observation process only. When aiming at picking a ‘good control’ there is thus a trade-off present since such controls must balance improving the estimation and minimizing the cost functional. The effect of considering strong formulations within stochastic control problems with unobservable components was recently reviewed in [14]. In particular, they illuminated the link between controlling the information flow and the notions of dual effect (referring to the control’s effect on higher order moments of the posterior distribution) and the much-studied trade-off between exploration and exploitation (referring to the knowledge of the unobservable component contra cost optimisation); we refer to that article for further details and references on the topic. Following [14], we notably introduce our problem by defining a class of pre-admissible controls and restricting, in turn, to the subset of controls which rely on the information provided by the observations only. Within the literature on optimal control under partial information, controls that are thus restricted to depend on the observation process only are often referred to as ‘strict-sense’ controls. This is in contrast to so-called ‘wide-sense’ controls which are allowed to depend on a somewhat larger filtration; see e.g. [9, 21, 24]. It is also common to define admissible controls on some reference probability space and let the controlled dynamics enter the problem via a change of measure (rendering the filtration independent of the chosen control); see e.g. [2, 10, 29]. Our main result verifies that for the problem at hand, such variants of weak formulations would agree in value.
Related literature
Bayesian search problems have been studied from a more applied perspective in e.g. [1, 18, 26, 39]. This literature aside, there are essentially three streams of research to which our problem relates:
Bayesian stopping problems: Motivated by aims similar to ours, closely related problems have been considered within an optimal stopping framework. For example, in [19], the authors let an unobservable state influence the drift of the signal process and their aim is to estimate this state as accurately but also as quickly as possible. Considering a Bayesian framework, they formulate it as an optimal stopping problem which they subsequently address. An extended version of such a stopping problem was studied in [20] to which we refer for further references.
Control problems with hidden state: Our formulation includes the problem of minimising, over admissible controls, an objective of the form [^1] Notably there is still a trade-off present when choosing the control since it affects both the cost and the estimation of the unknown signal; as before, this is a consequence of the fact that we consider a strong formulation where the choice of the control influences the available information through the filtration.
(4) $\mathbb{E}\int_{0}^{\infty}e^{-\beta t}\hat{k}(X,u_{t})\mathrm{d}t$
for some given cost function $\hat{k}$; this can be seen by defining $k(\mu,v)=\mu(\hat{k}(\cdot,v))$, for any probability measure $\mu$ and action $v$. Closely related problems have been studied previously in the literature: In [9] and [29], the authors let the observation process follow the dynamics corresponding to letting $h(v,x)=vx$ in (1), for real-valued controls, and the aim is to optimise an objective similar to (4) but with a running cost depending on the observation process itself; specifically, the cost takes the form $\hat{k}(Y^{u}_{t})$ for some function $\hat{k}$ satisfying certain properties. The corresponding finite-horizon problem was studied in [30]. More recently, the above-mentioned article [14], considered an observation process whose dynamics is again similar to (1) but with $h(u_{t},X)$ replaced by $b(t,Y^{u}_{t},u_{t})X$ for some function $b$ (they allowed for a multi-dimensional signal) and with a similar controlled volatility coefficient appearing in front of the noise term. They then optimised a finite-horizon objective with a running cost of the form $\hat{k}(t,Y^{u}_{t},X,u_{t})$. While these papers motivated our work, our setup is different: We crucially allow for a non-linear dependence on the signal in the drift of the observation process and a non-linear dependence on the posterior distribution in the cost functional. In order to render the presentation focused on the difficulties arising from this non-linearity, we develop the theory for the case where neither the drift nor the cost has any explicit dependence on the observation process.
Stochastic control under partial information: Our problem also bears resemblance to a large class of problems known as stochastic control under partial information. The difference between our setup and this class of problems is that we keep our signal constant while we control the observation rather than the signal itself. In comparison, in the partial observation literature, one is typically interested in a problem where the hidden signal, say $X_{t}$, $t\geq 0$, follows itself a controlled Itô dynamics while the observation is given by (1) — with $X$ replaced by $X_{t}$ — for some non-controlled [^2] Notably, [2], [31] and [38] also allow for controlled observations. function $h(x)$, $x\in\mathbb{R}$; the aim is to optimise an objective of the form (4) — with $X$ replaced by $X_{t}$. The nomenclature ‘partially observable’ refers to the fact that controls must be adapted to the filtration generated by the observation process; controls given in feedback form in terms of the signal are in general not admissible.
The predominant approach has been to address this problem via a reformulation enabling viewing it as a control problem where the conditional distribution of the hidden signal plays the role of the controlled state variable. In its original form, this approach amounts to rewriting the objective onto the form (2) where now $\pi^{u}_{t}=\mathrm{Law}(X_{t}|\mathcal{Y}^{u}_{t})$; this filter then solves the so-called Kushner-Stratonovich (KS) equation which is similar to (3) but with additional (controlled) terms appearing due to the dynamics of $X$ itself. At the cost of the controlled process taking values in the space of probability measures, the resulting problem of optimising (2) over solutions to the KS equation is conceptually simpler since it is a fully observable problem; for example, controls given in feedback form in terms of the filter are admissible. The strategy of linking the problem of original interest to such an observable control problem is often referred to as the separation principle. Provided the latter problem admits an optimiser, the hope is to construct a solution to the original problem too based on that solution. When $X_{0}$ is Gaussian and the various coefficients appearing in the dynamics of the signal and observation processes are linear, the resulting filter, referred to as the Kalman-Bucy filter, is Gaussian too and this problem can be solved explicitly (see e.g. [25], [10, Chapter 7] and [38, Section 7.5]). The study of the KS equation was also crucial for the analysis in [23] and [21] where the existence of optimal controls was established.
A more common line of attack, however, has been to employ the separation principle in conjunction with the unnormalised filter, which is a measure-valued process $\sigma_{t}$ satisfying $\pi^{u}_{t}(\varphi)=\sigma^{u}_{t}(\varphi)/\sigma^{u}_{t}(1)$, $\varphi\in C_{b}$. By considering a change of measure rendering the observation process a Brownian motion, one may express the cost functional in terms of this unnormalised filter and also obtain an equation for it, the so-called Zakai equation, which is a controlled measure-valued SDE driven by the observation process itself. One may again view the associated (separated) optimisation problem as a fully observable stochastic control problem featuring a controlled measure-valued process. Crucially, the Zakai equation turns out to be an easier equation in certain ways explaining the popularity of this method. For example, the study of the Zakai equation was a crucial component of the analysis in [24]. More pertinently, it is often assumed that the unnormalised conditional distribution admits a density: the Zakai equation then reduces to the so-called Duncan-Mortensen-Zakai equation for the density process itself. Within such a setup, there are numerous contributions to date: in [10, Chapter 8], the problem is dealt with by use of the stochastic maximum principle as well as via the DPP combined with semigroup arguments; in [28], it is proven that the value function is a viscosity solution (albeit in a rather weak sense) to an associated HJB equation; and in [34] and [27], it is assumed that the density process belongs to $L^{2}$ and an weighted $L^{2}$-space, respectively, and it is then shown that the value function is the unique viscosity solution to a corresponding HJB equation. Motivated by similar questions, [35] established an Itô formula for solutions to the KS and Zakai equations (see also [36] where viscosity solutions are discussed for some related equations).
In [2], the authors consider yet a different implementation of the separation principle. Introducing again the measure making the observations a Brownian motion, they consider the joint distribution under this measure, given the accumulated observations, of its (inverse) Radon-Nikodym derivative and the signal. The objective can clearly be rewritten in terms of this probability-measure–valued process. To find an equation for its dynamics, one acknowledges that the Radon-Nikodym derivative solves an SDE driven by the observations. Effectively, one is thus facing a situation analogous to the one of deriving the KS equation when the only thing being observed is one of the noise processes driving an ‘enlarged signal’; the associated normalised and unnormalised conditional probabilities then coincide and one easily obtains its dynamics. Again, one may then study the ‘separated problem’ associated with this objective and controlled probability-measure–valued dynamics. In essence, one has obtained a simpler dynamics at the cost of introducing an additional factor. The authors consider a non-Markovian version of this (separated) problem which they address by use of BSDE methods. In a closely related work, [3], the analogous problem is studied in a Markovian setup and it is shown that the value function is then the unique viscosity solution of an associated HJB equation formulated in terms of the so-called Lions derivative.
Finally, in [6], a closely related problem where one controls a signal while observing only one of its driving noise processes is studied. In their setup, however, the volatility coefficient in front of the observed noise term does not depend on the signal itself and therefore their formulation does not encompass more general partially observable problems as in [2, 3]. However, their approach is more closely related to ours in the sense that they too establish comparison working directly on the space of probability measures (as opposed to some lifted space). Our results are effectively orthogonal though since the terms appearing in their controlled measure dynamics are complementary to those appearing in (3).
Structure of the article
@@
The remainder of the article is organized as follows: In Section 2 we introduce our main problem of interest and express the dynamics of the involved processes in the observation filtration; in Section 3 we introduce our weak problem formulation as well as the piecewise constant approximate formulation and provide our main result; Section 4 presents our stability results and Section 5 is devoted to the proof of our main result.
@@
Throughout, $\mu_{0}$ will be a given fixed probability measure on $\mathbb{R}$ with compact support. We denote by $\mathcal{P}$ the space of probability measures on $\mathbb{R}$ whose support is contained in the support of $\mu_{0}$; we equip this space with the topology of weak convergence rendering it a Polish space. Since $\mu_{0}$ has compact support, we can work with e.g. the Wasserstein-$1$ metric, which we denote by $\mathcal{W}$. That $\mu_{0}$ has compact support also implies that the space $\mathcal{P}$ is compact. We denote by $\mathcal{P}^{s}$ the (closed) subset of probability measures supported in one single point. We write $C_{b}$ for the bounded continuous functions on $\mathbb{R}$. Finally, for any $\mu\in\mathcal{P}$ and $\varphi:\mathbb{R}\to\mathbb{R}$, we set $\mu(\varphi):=\int_{\mathbb{R}}\varphi(x)\mu(\mathrm{d}x)$.